OpenStudy (lxelle):

Integrate x^1/2.lnx dx

OpenStudy (alexandervonhumboldt2):

@Directrix @perl @ganeshie8 @jim_thompson5910 htye should know

OpenStudy (alexandervonhumboldt2):

@wio

OpenStudy (shrutipande9):

let u=ln x thus du/dx =1/x ie du=(1/x) dx and then can u try to solve further?

OpenStudy (lxelle):

is lnx consider as logarithmic or exponential?

OpenStudy (shrutipande9):

i considered log here... @wio is way better than me...can u help?

OpenStudy (lxelle):

dv/dx = x^1/2 so, v = 2/3 x^3/2 right?

OpenStudy (shrutipande9):

yep..:)

OpenStudy (anonymous):

Is it: \[ \int x^{1/2}\ln (x)~dx \]

OpenStudy (perl):

you can do integration by parts

OpenStudy (lxelle):

Thanks @shrutipande9. Yes, @wio. Noted, @perl

jimthompson5910 (jim_thompson5910):

You'll use integration by parts \[\large \int x^{1/2}*\ln(x)dx\] \[\large \int \ln(x)*x^{1/2}dx\] u = ln(x), du = (1/x)dx dv = x^(1/2)dx, v = (2/3)x^(3/2) \[\Large \int u dv = uv - \int v du\] \[\large \int \ln(x)*x^{1/2}dx = \ln(x)*\left(\frac{2}{3}x^{3/2}\right)-\int \frac{2}{3}x^{3/2}\frac{1}{x}dx\] I'll let you finish

OpenStudy (lxelle):

oops, im stuck. D: @jim_thompson5910

OpenStudy (lxelle):

I got 2/3 x ^3/2 lnx - 2/3 (x^1/2)

jimthompson5910 (jim_thompson5910):

u = ln(x), du = dx/x dv = x^(1/2)dx, v = (2/3)x^(3/2) \[\Large \int u dv = uv - \int v du\] \[\large \int \ln(x)*x^{1/2}dx = \ln(x)*\left(\frac{2}{3}x^{3/2}\right)-\int \frac{2}{3}x^{3/2}\frac{1}{x}dx\] \[\large \int \ln(x)*x^{1/2}dx = \ln(x)*\left(\frac{2}{3}x^{3/2}\right)-\int \frac{2}{3}x^{3/2}x^{-1}dx\] \[\large \int \ln(x)*x^{1/2}dx = \ln(x)*\left(\frac{2}{3}x^{3/2}\right)-\frac{2}{3}\int x^{3/2-1}dx\] \[\large \int \ln(x)*x^{1/2}dx = \ln(x)*\left(\frac{2}{3}x^{3/2}\right)-\frac{2}{3}\int x^{1/2}dx\] How about now?

OpenStudy (lxelle):

exactly what i replied you before.

jimthompson5910 (jim_thompson5910):

I see. You left off the integral sign, but that's ok

OpenStudy (shrutipande9):

@jim_thompson5910 i am not that good in int...:/ was going right? sorry if i misleaded u @LXelle ..i tried..:)

jimthompson5910 (jim_thompson5910):

so what is \[\Large \int x^{1/2} dx\] equal to?

OpenStudy (lxelle):

@shrutipande9 It's okay. :)) You're doing fine. @jim_thompson5910 2/3 x ^3/2 right?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (lxelle):

THANKS ALL! YOU GUYS ARE AMAZING! :)

jimthompson5910 (jim_thompson5910):

u = ln(x), du = dx/x dv = x^(1/2)dx, v = (2/3)x^(3/2) \[\Large \int u dv = uv - \int v du\] \[\large \int \ln(x)*x^{1/2}dx = \ln(x)*\left(\frac{2}{3}x^{3/2}\right)-\int \frac{2}{3}x^{3/2}\frac{1}{x}dx\] \[\large \int \ln(x)*x^{1/2}dx = \ln(x)*\left(\frac{2}{3}x^{3/2}\right)-\int \frac{2}{3}x^{3/2}x^{-1}dx\] \[\large \int \ln(x)*x^{1/2}dx = \ln(x)*\left(\frac{2}{3}x^{3/2}\right)-\frac{2}{3}\int x^{3/2-1}dx\] \[\large \int \ln(x)*x^{1/2}dx = \ln(x)*\left(\frac{2}{3}x^{3/2}\right)-\frac{2}{3}\int x^{1/2}dx\] \[\large \int \ln(x)*x^{1/2}dx = \ln(x)*\left(\frac{2}{3}x^{3/2}\right)-\frac{2}{3}*\left(\frac{2}{3}x^{3/2}\right)+C\] \[\large \int \ln(x)*x^{1/2}dx = \frac{2}{3}x^{3/2}*\ln(x)-\frac{4}{9}*x^{3/2}+C\] Optionally you can factor out (2/3)x^(3/2) and you'd get \[\large \int \ln(x)*x^{1/2}dx = \frac{2}{3}x^{3/2}\left(\ln(x)-\frac{2}{3}\right)+C\]